OpenStudy (warmjen):

Find the present value, using the present value formula and a calculator. Achieve $225,500 at 8.65% compounded continuously for 8 years, 135 days.

OpenStudy (wolf1728):

1 + continuous rate = e^(annual rate) where 'e' is the mathematical constant 2.718281828 1 + continuous rate = 2.718281828^.0865 1 + continuous rate = 1.0903513677 continuous rate = .0903513677 or 9.03513677% Total = princ*(1.0903513677)^8.3698630137 Principal = Total / (1.0903513677)^8.3698630137 Principal = 225,500 / 2.0626532728 Principal = $109,325.21

OpenStudy (warmjen):

thank you so much :)